Operation Summary for November 8, 2010 - November 10, 2010 |
Monday Nov. 8 The day shift began with Operations monitoring Tevatron (TeV) store 8250 (148.49E30) and the Recycler stash (143.3E10), with the Antiproton Source (Pbar) stacking (8.5E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-1004 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam. During the day shift, FESS personnel continued their efforts to repair boiler #2 at CUB. During the evening shift, Operations terminated store 8250 at 6 PM, put the TeV into HEP shot setup, and established store 8251 at 7:30 PM with an average initial luminosity (AIL) of 314E30. Operations reported a short amount of downtime due to a Linac Klystron RF station (KRF3), a Linac quadrupole power supply (L:QPS104) that tripped off due to an I- Source trip, and a LRF2 trip. Tuesday Nov. 9 The midnight shift began with Operations monitoring store 8251 (153.82E30) and the stash (133.9E10), with Pbar stacking (22.9E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators tuned the Proton Source. Operations reported good stable running prior to the planned TeV A-E access to inspect and repair the TeV sector D11 power lead. During the day shift, Operations terminated store 8251 at 8:20 AM. Operations prepared CDF, D0, and the TeV A-E areas for access and necessary repairs. Operations reported that a problem with a D0 MVA gate interlock micro-switch caused the C-sector interlocks to drop out multiple times. Interlock technicians repaired the switch. Operations reported that a Pbar magnet (PMag) developed a ground fault at 10:30 AM. Replacement of the magnet began shortly after noon and continued through to the end of the shift. Water techs accessed the Pbar rings to repair LCW leaks. Operations had the TeV ramping by 2:35 PM. Operations put the TeV through dry squeezes and wet squeezes. TeV personnel checked the TeV orbits. MI personnel conducted beam studies. During the evening shift, TeV personnel completed their orbit checks. Operations put the TeV into shot setup at 5:45 PM. Operations reported that problems with C2 BPM flash updates and coalescing delayed the shot setup. Operations established store 8255 at 10:05 PM with an AIL of 340E30. Operations had the Pbar rings back on and in standby by 7:30 PM. The PMag replacement continued. Wednesday Nov. 10 The midnight shift began with Operations monitoring store 8255 (160.8E30) and the stash (7.2E10), with Pbar off due to a magnet change, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that the PMag replacement had been completed by 4 AM; Operations reestablished stacking beam at 4:45 AM. Operations reported that the E1 wet engine had tripped off on multiple occasions. The Plans for Wednesday The plans for today are to stack, stash, and store. The Run Coordinator will allow cryo system personnel to work on the E1 wet engine, for MI personnel to conduct studies, for Linac work, and for a Booster pump switch due to a leaking seal. |
|
More Information Comments and Suggestions |
| last modified 11/10/2010 email Fermilab |
| Security, Privacy, Legal | |